### v18

# Mathematics with the knowledge predicate ${\mathcal{K}}$ satisfied by every currently known theorem Apoloniusz Tyszka

Abstract
${\mathcal{K}}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known theorems. The set ${\mathcal{K}}$ is time-dependent, publicly available, and contains theorems both from formal and constructive mathematics. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}$. Mathematical statements with known constructive proofs exist in ${\mathcal{K}}$ separately and form the set ${\mathcal{K}}_{\textstyle c}\subseteq{\mathcal{K}}$. We assume that mathematical sets are atemporal entities. They exist formally in $ZFC$ theory although their properties can be time-dependent (when they depend on ${\mathcal{K}}$) or informal. Algorithms always terminate. We explain the distinction between algorithms whose existence is provable in $ZFC$ and constructively defined algorithms which are currently known. By using this distinction, we obtain non-trivially true statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ that belong to constructive and informal mathematics and refer to the current mathematical knowledge on ${\mathcal{X}}$. We present a non-trivially true statement in computability (which may be falsified) in which the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge appears exclusively inside the statement.
Key words and phrases: constructive mathematics, constructively defined algorithms, current mathematical knowledge, informal mathematics, known algorithms, time-dependent notion of truth in constructive mathematics.
2020 Mathematics Subject Classification: 03F65.
1 Introduction
This article in an extended and changed version of [24].
${\mathcal{K}}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known theorems. The set ${\mathcal{K}}$ is time-dependent, publicly available, and contains theorems both from formal and constructive mathematics. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}$. Mathematical statements with known constructive proofs exist in ${\mathcal{K}}$ separately and form the set ${\mathcal{K}}_{\textstyle c}\subseteq{\mathcal{K}}$. We assume that mathematical sets are atemporal entities. They exist formally in $ZFC$ theory although their properties can be time-dependent (when they depend on ${\mathcal{K}}$) or informal. The true statement "${\mathcal{K}}$ is non-empty" is outside ${\mathcal{K}}$ forever because ${\mathcal{K}}$ is not a formal set.
Paul Cohen proved in $1963$ that the equality $2^{\aleph_{0}}=\aleph_{1}$ is independent of $ZFC$, see [3]. Before $1963$, the statement "There is a constructively defined integer $n\geqslant 1$ such that $2^{\aleph_{0}}=\aleph_{n}$" was outside ${\mathcal{K}}$. Since $1963$, this statement is outside ${\mathcal{K}}$ forever. The true statement "There exists a set ${\mathcal{X}}\subseteq\{1,\ldots,49\}$ such that ${\rm card}({\mathcal{X}})=6$ and ${\mathcal{X}}$ never occurred as the winning six numbers in the Polish Lotto lottery" refers to the current non-mathematical knowledge and is outside ${\mathcal{K}}$ forever.
Algorithms always terminate. Semi-algorithms may not terminate. There is the distinction between existing algorithms (i.e. algorithms whose existence is provable in $ZFC$) and known algorithms (i.e. algorithms whose definition is constructive and currently known), see [2], [17, p. 9]. By using this distinction, we obtain non-trivially true statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ that belong to constructive and informal mathematics and refer to the current mathematical knowledge on ${\mathcal{X}}$. For every such statement $\Phi$, Observations 14 hold.
Observation 1.
The truth of $\Phi$ concerning the current mathematical knowledge is implied by a true statement of the form:
(the conjunction of $i$ conditions of the form $\alpha\in{\mathcal{K}}$) $\wedge$ (the conjunction of $j$ conditions of the form $\beta\not\in{\mathcal{K}}$) $\wedge$
(the conjunction of $k$ conditions of the form $\gamma\in{\mathcal{K}}_{\textstyle c}$) $\wedge$ (the conjunction of $l$ conditions of the form $\delta\not\in{\mathcal{K}}_{\textstyle c}$),
where $i,j,k,l\in\mathbb{N}$ and $\alpha,\beta,\gamma,\delta$ are mathematical statements. In particular, the truth of Statement 6 concerning the current mathematical knowledge guarantees that the set
${\mathcal{X}}=\{n\in\mathbb{N}:~{}the~{}interval~{}[-1,n]~{}contains~{}more~{}than~{}29.5+\frac{\textstyle 11!}{\textstyle 3n+1}\cdot\sin(n)~{}primes~{}of~{}the~{}form~{}k!+1\}$
satisfies the following conjunction:
${\rm(the~{}statement~{}}"the~{}algorithm~{}{\mathcal{A}}_{1}~{}returns~{}the~{}logical~{}value~{}of~{}the~{}statement~{}{\rm card}({\mathcal{X}})=\omega"{\rm~{}is~{}outside~{}{\mathcal{K}})\wedge\ldots\wedge}$
${\rm(the~{}statement~{}}"the~{}algorithm~{}{\mathcal{A}}_{p}~{}returns~{}the~{}logical~{}value~{}of~{}the~{}statement~{}{card}({\mathcal{X}})=\omega"{\rm~{}is~{}outside~{}{\mathcal{K}}),}$
where ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{p}$ are the all known algorithms with no input. Statement 6 will be false when someone proves that there is a constructively defined algorithm ${\mathcal{A}}$ with no input that returns the logical value of the statement ${\rm card}({\mathcal{X}})=\omega$.
Observation 2.
The proof of $\Phi$ uses mathematical theorems. For example, the proof of Statement 6 uses the fact that the set ${\mathcal{X}}$ from Observation 1 satisfies ${\rm card}({\mathcal{X}})<\omega\Rightarrow{\mathcal{X}}\subseteq(-\infty,501893]$.
Observation 3.
The proof of $\Phi$ does not use the assumption that we can list the all elements of ${\mathcal{K}}$.
Observation 4.
The proof of $\Phi$ does not use the assumption that we can list the all elements of ${\mathcal{K}}_{\textstyle c}$.
Observation 5 is known from the beginning of computability theory and shows that the predicate ${\mathcal{K}}$ increases intuitive mathematics.
Observation 5.
Church’s thesis is based on the fact that the all currently known intuitively computable functions are computable.
In Observation 6, the predicate ${\mathcal{K}}$ trivially increases constructive mathematics.
Observation 6.
The largest known prime number has the form $12^{\textstyle n}-$ 1.
2 Time-dependent notion of truth in constructive mathematics
Below is the English summary of [22] available at the internet address of [22].
The basic philosophical idea of intuitionism is that mathematical entities exist only as mental constructions and that the notion of truth of a proposition should be equated with its verification or the existence of proof. However different intuitionists explained the existence of a proof in fundamentally different ways. There seem to be two main alternatives: the actual and potential existence of a proof. The second proposal is also understood in two alternative ways: as knowledge of a method of construction of a proof or as knowledge-independent and tenseless existence of a proof. This paper is a presentation and analysis of these alternatives.
In constructive mathematics ([16]) and the traditional Brouwerian intuitionism ([11, p. 135]), the truth of a mathematical statement means that we know a constructive proof. Therefore, the truth of a mathematical statement depends on time, where the statement is formally stated in the classical mathematics without the predicate ${\mathcal{K}}$.
In this article, mathematical statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ refer to time because they refer to the current mathematical knowledge on ${\mathcal{X}}$. They cannot be formally stated in the classical mathematics without the predicate ${\mathcal{K}}$ and their logical values may change in time.
In Martin-Löf’s terminology ([10, p. 142]), every currently known theorem is actually true whereas every theorem (known or unknown) is potentially true. Actual truth is knowledge dependent and tensed. Potential truth is knowledge independent and tenseless.
3 Basic definitions and examples
Definition 1 applies to sets ${\mathcal{X}}\subseteq\mathbb{N}$ whose infiniteness is false or unproven.
Definition 1.
[…]
The set of known elements of ${\mathcal{X}}$ is finite and time-dependent, so cannot be defined in the formal language of classical mathematics. Let $t$ denote the largest twin prime that is smaller than ((((((((9!)!)!)!)!)!)!)!)!. The number $t$ is an unknown element of the set of twin primes.
Definition 2.
Conditions (1)-(5) concern sets ${\mathcal{X}}\subseteq\mathbb{N}$. (1) A known algorithm with no input returns an integer $n$ satisfying ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\mathcal{X}}\subseteq(-\infty,n]$. (2) A known algorithm for every $k\in\mathbb{N}$ decides whether or not $k\in{\mathcal{X}}$. (3) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}({\mathcal{X}})=\omega$" is outside ${\mathcal{K}}$. (4) There are many elements of ${\mathcal{X}}$ and it is conjectured, though so far unproven, that ${\mathcal{X}}$ is infinite. (5) ${\mathcal{X}}$ is naturally defined. The infiniteness of ${\mathcal{X}}$ is false or unproven. ${\mathcal{X}}$ has the simplest definition among known sets ${\mathcal{Y}}\subseteq\mathbb{N}$ with the same set of known elements.
Condition (3) implies that no known proof shows the finiteness/infiniteness of ${\mathcal{X}}$. No known set ${\mathcal{X}}\subseteq\mathbb{N}$ satisfies Conditions (1)-(4) and is widely known in number theory or naturally defined, where this term has only informal meaning.
Let []delimited-[][$\cdot$][ ⋅ ] denote the integer part function.
Example 1.
The set
${\mathcal{X}}=\left\{\begin{array}[]{cl}\mathbb{N},&{\rm if~{}}[\frac{((((((((9!)!)!)!)!)!)!)!)!}{\pi}]~{}is~{}odd\\ \emptyset,&otherwise\end{array}\right.$
does not satisfy Condition (3) because we know an algorithm with no input that computes $[\frac{((((((((9!)!)!)!)!)!)!)!)!}{\pi}]$. The set of known elements of ${\mathcal{X}}$ is empty. Hence, Condition (5) fails for ${\mathcal{X}}$.
Example 2.
([2][17, p. 9]). The function
$\mathbb{N}\ni n\stackrel{{\scriptstyle\textstyle h}}{{\longrightarrow}}\left\{\begin{array}[]{rl}1,&if~{}the~{}decimal~{}expansion~{}of~{}\pi~{}contains~{}n~{}consecutive~{}zeros\\ 0,&otherwise\end{array}\right.$
is computable because $h=\mathbb{N}\times\{1\}$ or there exists $k\in\mathbb{N}$ such that
$h=(\{0,\ldots,k\}\times\{1\})\cup(\{k+1,~{}k+2,~{}k+3,~{}\ldots\}\times\{0\})$
No known algorithm computes the function $h$.
Example 3.
The set
${\mathcal{X}}=\left\{\begin{array}[]{rl}\mathbb{N},&if~{}the~{}continuum~{}hypothesis~{}holds\\ \emptyset,&otherwise\end{array}\right.$
is decidable. This ${\mathcal{X}}$ satisfies Conditions (1) and (3) and does not satisfy Conditions (2), (4), and (5). These facts will hold forever.
4 A consequence of the physical limits of computation
Statement 1.
No set ${\mathcal{X}}\subseteq\mathbb{N}$ will satisfy Conditions (1)-(4) forever, if for every algorithm with no input, at some future day, a computer will be able to execute this algorithm in $1$ second or less.
Proof.
The proof goes by contradiction. We fix an integer $n$ that satisfies Condition (1). Since Conditions (1)-(3) will hold forever, the semi-algorithm in Figure 1 never terminates and sequentially prints the following sentences:
$n+1\not\in{\mathcal{X}},~{}n+2\not\in{\mathcal{X}},~{}n+3\not\in{\mathcal{X}},~{}\ldots$ (T)
Figure 1.   Semi-algorithm that terminates if and only if ${\mathcal{X}}$ is infinite.
The sentences from the sequence (T) and our assumption imply that for every integer $m>n$ computed by a known algorithm, at some future day, a computer will be able to confirm in $1$ second or less that $(n,m]\cap{\mathcal{X}}=\emptyset$. Thus, at some future day, numerical evidence will support the conjecture that the set ${\mathcal{X}}$ is finite, contrary to the conjecture in Condition (4). ∎
The physical limits of computation ([9]) disprove the assumption of Statement 1.
5 Statements which refer to Conditions (1)-(5)
Edmund Landau’s conjecture states that the set ${\mathcal{P}}_{n^{2}+1}$ of primes of the form $n^{2}+1$ is infinite, see [20], [21], [27].
Statement 2.
Condition (1) remains unproven for ${\mathcal{X}}={\mathcal{P}}_{n^{2}+1}$.
Proof.
For every set ${\mathcal{X}}\subseteq\mathbb{N}$, there exists an algorithm ${\tt Alg}({\mathcal{X}})$ with no input that returns
$n=\left\{\begin{array}[]{rl}0,&{\rm if~{}}{\rm card}({\mathcal{X}})\in\{0,\omega\}\\ \max({\mathcal{X}}),&{\rm otherwise}\end{array}\right.$
This $n$ satisfies the implication in Condition (1), but the algorithm ${\tt Alg}({\mathcal{P}}_{n^{2}+1})$ is unknown because its definition is ineffective. ∎
Statement 3.
The statement
$\exists n\in\mathbb{N}~{}({\rm card}({\mathcal{P}}_{n^{2}+1})<\omega\Rightarrow{\mathcal{P}}_{n^{2}+1}\subseteq[2,n+3])$
remains unproven in $ZFC$ and classical logic without the law of excluded middle.
Statement 4.
The set
${\mathcal{X}}=\{k\in\mathbb{N}:~{}{\rm card}({\mathcal{P}}_{n^{2}+1}\cap[-1,k])>10^{\textstyle 10^{\textstyle 10}}\}\cup\{k\in{\mathcal{P}}_{n^{2}+1}:~{}{\rm card}({\mathcal{P}}_{n^{2}+1}\cap[-1,k])\leqslant 10^{\textstyle 10^{\textstyle 10}}\}$
satisfies Conditions (2)-(4). Condition (1) fails for ${\mathcal{X}}$, ${\rm card}({\mathcal{X}})<\omega\Rightarrow{\rm card}({\mathcal{X}})\leqslant 10^{\textstyle 10^{\textstyle 10}}$.
Proof.
Since ${\rm card}({\mathcal{P}}_{n^{2}+1}\cap[2,10^{28}))=2199894223892$ ([21]) and the inequality ${\rm card}(P_{n^{2}+1})\geqslant$ $10^{\textstyle 10^{\textstyle 10}}$ remains unproven, Conditions (3) and (4) hold. ∎
For a non-negative integer $n$, let $\theta(n)$ denote the largest integer divisor of $10^{\textstyle 10^{\textstyle 10}}$ smaller than $n$. Let $\kappa:\mathbb{N}\to\mathbb{N}$ be defined by setting $\kappa(n)$ to be the exponent of $2$ in the prime factorization of $n+1$.
Statement 5.
([25, p. 250]). The set
${\mathcal{X}}=\{n\in\mathbb{N}:(\theta(n)+\kappa(n))^{2}+1~{}is~{}prime\}$
satisfies Conditions (1)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined. Condition (1) holds with $n=10^{\textstyle 10^{\textstyle 10}}$.
Let ${\mathcal{P}}_{n!+1}$ denote the set of primes of the form $n!+1$.
Conjecture 1.
([1, p. 443], [6]). The set ${\mathcal{P}}_{n!+1}$ is infinite.
For a non-negative integer $n$, let $\rho(n)$ denote $29.5+\frac{\textstyle 11!}{\textstyle 3n+1}\cdot\sin(n)$.
Statement 6.
The set
${\mathcal{X}}=\{n\in\mathbb{N}:the~{}interval~{}[-1,n]~{}contains~{}more~{}than~{}\rho(n)~{}elements~{}of~{}{\mathcal{P}}_{n!+1}\}$
satisfies Conditions (1)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined. $501893\in{\mathcal{X}}$. Condition (1) holds with $n=501893$. ${\rm card}({\mathcal{X}}\cap[0,501893])=159827$. ${\mathcal{X}}\cap[501894,\infty)=$$\{n\in\mathbb{N}:$ $the$ $interval$ $[-1,n]$ $contains$ $at$ $least$ $30$ $elements$ $of$ ${\mathcal{P}}_{n!+1}\}$.
Proof.
For every integer $n\geqslant 11!$, $30$ is the smallest integer greater than $\rho(n)$. By this, if $n\in{\mathcal{X}}\cap[11!,\infty)$, then $n+1,~{}n+2,~{}n+3,~{}\ldots~{}\in{\mathcal{X}}$. Hence, Condition (1) holds with $n=11!-1$. Since the inequality ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 30$ remains unproven, Condition (3) holds. The interval $[-1,11!-1]$ contains exactly three primes of the form $k!+1$: $1!+1$, $2!+1$, $3!+1$. For every integer $n>503000$, the inequality $\rho(n)>3$ holds. Therefore, the execution of the following MuPAD code
m:=0:
for n from 0.0 to 503000.0 do
if n<1!+1 then r:=0 end_if:
if n>=1!+1 and n<2!+1 then r:=1 end_if:
if n>=2!+1 and n<3!+1 then r:=2 end_if:
if n>=3!+1 then r:=3 end_if:
if r>29.5+(11!/(3*n+1))*sin(n) then
m:=m+1:
print([n,m]):
end_if:
end_for:
displays the all known elements of ${\mathcal{X}}$. The output ends with the line $[501893.0,~{}159827]$, which proves Condition (1) with $n=501893$ and Condition (4) with ${\rm card}({\mathcal{X}})\geqslant 159827$. ∎
T. Nagell proved in [14] (cf. [19, p. 104]) that the equation $x^{2}-17=y^{3}$ has exactly $16$ integer solutions, namely (±$(\pm 3,-2)$, (±$(\pm 4,-1)$, (±$(\pm 5,2)$, (±$(\pm 9,4)$, (±$(\pm 23,8)$, (±$(\pm 282,43)$, (±$(\pm 375,52)$, (±$(\pm 378661,5234)$. The set
$\bigcup_{\stackrel{{\scriptstyle\textstyle(x,y)\in\mathbb{Z}\times\mathbb{Z}}}{{\textstyle(x^{2}-y^{3}-17)\cdot(y^{2}-x^{3}-17)=0}}}\{(x+8)^{8}\}$
has exactly $23$ elements. Among them, there are $14$ integers from the interval $[1,2199894223892]$. Let ${\mathcal{W}}$ denote the set
$\bigcup_{\stackrel{{\scriptstyle\textstyle(x,y)\in\mathbb{Z}\times\mathbb{Z}}}{{\textstyle(x^{2}-y^{3}-17)\cdot(y^{2}-x^{3}-17)=0}}}\{k\in\mathbb{N}:~{}k~{}is~{}the~{}(x+8)^{8}-th~{}element~{}of~{}{\mathcal{P}}_{n^{2}+1}\}$
From [21], it is known that ${\rm card}({\mathcal{P}}_{n^{2}+1}\cap[2,10^{28}))=2199894223892$. Hence, ${\rm card}({\mathcal{W}}\cap[2,10^{28}))=14$ and $14$ elements of ${\mathcal{W}}$ can be practically computed. The inequality ${\rm card}({\mathcal{P}}_{n^{2}+1})\geqslant(378661+8)^{8}$ remains unproven. The last two sentences and Statement 6 imply the following corollary.
Corollary 1.
If we add ${\mathcal{W}}$ to ${\mathcal{X}}$, then the following statements hold: Condition (1) fails for ${\mathcal{X}}$, $159827+14\leqslant{\rm card}({\mathcal{X}})$, the above lower bound is currently the best known, ${\rm card}({\mathcal{X}})<\omega\Rightarrow{\rm card}({\mathcal{X}})\leqslant 159827+23$, the above upper bound is currently the best known, ${\mathcal{X}}$ satisfies Conditions (2)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined.
Corollary 2.
Since the inequality ${\rm card}({\mathcal{P}}_{n^{2}+1})>9^{\textstyle 999}$ remains unproven and $10^{\textstyle 953}<9^{\textstyle 999}<10^{\textstyle 954}$, analogical statements hold when we add to ${\mathcal{X}}$ the set
$\bigcup_{\textstyle i\in\mathbb{N}}\left\{k\in\mathbb{N}:~{}k-501894~{}is~{}the~{}\left(\left[\frac{9^{\textstyle 999}}{10^{\textstyle i}}\right]+1\right)-th~{}element~{}of~{}{\mathcal{P}}_{n^{2}+1}\right\}$
which has at most $955$ elements.
For a non-negative integer $i$, let $d(i)$ denote the smallest prime divisor of $\left[31+\frac{\textstyle 10^{\textstyle 6}}{\textstyle i+1}\right]$.
Statement 7.
The set
${\mathcal{X}}=\bigcup_{\textstyle i\in\mathbb{N}}\left\{k^{\textstyle i}:~{}k~{}is~{}the~{}d(i)-th~{}element~{}of~{}{\mathcal{P}}_{n!+1}\right\}$
satisfies Conditions (2)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined. Condition (1) fails for ${\mathcal{X}}$. ${\rm card}({\mathcal{X}})\geqslant 946732$. If ${\rm card}({\mathcal{P}}_{n!+1})\leqslant 28$, then ${\rm card}({\mathcal{X}})=946732$. If $29\leqslant{\rm card}({\mathcal{P}}_{n!+1})\leqslant 30$, then ${\rm card}({\mathcal{X}})=946745$. If ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 31$, then ${\rm card}({\mathcal{X}})=\omega$.
Proof.
The inequality ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 23$ holds, see [5]. The inequality ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 29$ remains unproven. The execution of the following MuPAD code
[m,n]:=[0,0]:
for i from 0 to 10^6-1 do
A:=numlib::primedivisors(floor(31+(10^6/(i+1)))):
if A[1]<=23 then m:=m+1 end_if:
if A[1]<=29 then n:=n+1 end_if:
end_for:
print([m,n]):
displays $[946732,946745]$. The last claim of Statement 7 holds because $d\left(\left[31+\frac{\textstyle 10^{\textstyle 6}}{\textstyle i+1}\right]\right)=31$ for every integer $i\geqslant 10^{\textstyle 6}$. Condition (1) fails for ${\mathcal{X}}$ because we cannot rule out the possibility that $29\leqslant{\rm card}({\mathcal{P}}_{n!+1})\leqslant 30$. ∎
6 Satisfiable conjunctions which consist of Conditions (1)-(5) and their negations
Open Problem 1.
Is there a set ${\mathcal{X}}\subseteq\mathbb{N}$ which satisfies Conditions (1)-(5)?
Open Problem 1 asks about the existence of a year $t\geqslant 2024$ in which the conjunction
$(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})$
will hold for some ${\mathcal{X}}\subseteq\mathbb{N}$. For every year $t\geqslant 2024$ and for every $i\in\{1,2,3\}$, a positive solution to Open Problem $i$ in the year $t$ may change in the future. Currently, the answers to Open Problems 15 are negative.
The set ${\mathcal{X}}={\mathcal{P}}_{n^{2}+1}$ satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})$
The set ${\mathcal{X}}=\{0,\ldots,10^{6}\}\cup{\mathcal{P}}_{n^{2}+1}$ satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge\neg(Condition~{}{\tt 5})$
Let $f(1)=10^{6}$, and let $f(n+1)=f(n)^{f(n)}$ for every positive integer $n$. The set
${\mathcal{X}}=\left\{\begin{array}[]{l}\mathbb{N},~{}if~{}2^{\textstyle 2^{\textstyle f(9^{9})}}+1~{}is~{}composite\\ \{0,\ldots,10^{6}\},~{}otherwise\end{array}\right.$
satisfies the conjunction
$(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge\neg(Condition~{}{\tt 5})$
Open Problem 2.
Is there a set ${\mathcal{X}}\subseteq\mathbb{N}$ that satisfies the conjunction
$(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})?$
The numbers $12^{\textstyle 2^{k}}+$ 1 are prime for $k\in\{0,1,2,3,4\}$. It is open whether or not there are infinitely many primes of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 158] and [15, p. 74]. It is open whether or not there are infinitely many composite numbers of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 159] and [15, p. 74]. Most mathematicians believe that $12^{\textstyle 2^{k}}+$ 1 is composite for every integer $k\geqslant 5$, see [7, p. 23]. The set
${\mathcal{X}}=\left\{\begin{array}[]{l}\mathbb{N},~{}if~{}2^{\textstyle 2^{\textstyle f(9^{9})}}+1~{}is~{}composite\\ \{0,\ldots,10^{6}\}\cup\\ \{n\in\mathbb{N}:n~{}is~{}the~{}sixth~{}prime~{}number~{}of~{}the~{}form~{}2^{\textstyle 2^{\textstyle k}}+1\},~{}otherwise\end{array}\right.$
satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge\neg(Condition~{}{\tt 5})$
Open Problem 3.
Is there a set ${\mathcal{X}}\subseteq\mathbb{N}$ that satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})?$
It is possible, although very doubtful, that at some future day, the set ${\mathcal{X}}={\mathcal{P}}_{n^{2}+1}$ will solve Open Problem 2. The same is true for Open Problem 3. It is possible, although very doubtful, that at some future day, the set ${\mathcal{X}}=\{k\in\mathbb{N}:2^{\textstyle 2^{k}}+1~{}is~{}composite\}$ will solve Open Problem 1. The same is true for Open Problems 2 and 3.
Table 1 shows satisfiable conjunctions of the form
#$\#{\rm(Condition}~{}{\tt 1}{\rm)}\wedge{\rm(Condition}~{}{\tt 2}{\rm)}\wedge\#{\rm(Condition}~{}{\tt 3}{\rm)}\wedge{\rm(Condition}~{}{\tt 4}{\rm)}\wedge\#{\rm(Condition}~{}{\tt 5}{\rm)}$
where ##\## denotes the negation $\neg$ or the absence of any symbol.
Table 1.   Five satisfiable conjunctions.
7 Statements which refer to Conditions (1a)-(5a) and (6)-(11)
Definition 3.
Conditions (1a)-(5a) concern sets ${\mathcal{X}}\subseteq\mathbb{N}$. (1a) A known algorithm with no input returns a positive integer $n$ satisfying ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\mathcal{X}}\subseteq(-\infty,n]$. (2a) A known algorithm for every $k\in\mathbb{N}$ decides whether or not $k\in{\mathcal{X}}$. (3a) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}({\mathcal{X}})<\omega$" is outside ${\mathcal{K}}$. (4a) There are many elements of ${\mathcal{X}}$ and it is conjectured, though so far unproven, that ${\mathcal{X}}$ is finite. (5a) ${\mathcal{X}}$ is naturally defined. The finiteness of ${\mathcal{X}}$ is false or unproven. ${\mathcal{X}}$ has the simplest definition among known sets ${\mathcal{Y}}\subseteq\mathbb{N}$ with the same set of known elements.
Statement 8.
The set
${\mathcal{X}}=\Bigl{\{}n\in\mathbb{N}:~{}the~{}interval~{}[-1,n]~{}contains~{}more~{}than~{}6.5+\frac{10^{6}}{3n+1}\cdot\sin(n)~{}squares~{}of~{}the~{}form~{}k!+1\Bigr{\}}$
satisfies Conditions (1a)-(5a) except the requirement that ${\mathcal{X}}$ is naturally defined. $95151\in{\mathcal{X}}$. Condition (1a) holds with $n=95151$. ${\rm card}({\mathcal{X}}\cap[0,95151])=30311$. ${\mathcal{X}}\cap[95152,\infty)=$$\{n\in\mathbb{N}:$ $the$ $interval$ $[-1,n]$ $contains$ $at$ $least$ $7$ $squares$ $of$ $the$ $form$ $k!+1\}$.
Proof.
For every integer $n>10^{6}$, $7$ is the smallest integer greater than $6.5+\frac{10^{6}}{3n+1}\cdot\sin(n)$. By this, if $n\in{\mathcal{X}}\cap(10^{6},\infty)$, then $n+1,~{}n+2,~{}n+3,~{}\ldots~{}\in{\mathcal{X}}$. Hence, Condition (1a) holds with $n=10^{6}$. It is conjectured that $k!+1$ is a square only for $k\in\{4,5,7\}$, see [26, p. 297]. Hence, the inequality ${\rm card}(\{k\in\mathbb{N}\setminus\{0\}:k!+1~{}is~{}a~{}square\})>3$ remains unproven. Since $3<7$, Condition (3a) holds. The interval $[-1,10^{6}]$ contains exactly three squares of the form $k!+1$: $4!+1$, $5!+1$, $7!+1$. Therefore, the execution of the following MuPAD code
m:=0:
for n from 0.0 to 1000000.0 do
if n<25 then r:=0 end_if:
if n>=25 and n<121 then r:=1 end_if:
if n>=121 and n<5041 then r:=2 end_if:
if n>=5041 then r:=3 end_if:
if r>6.5+(1000000/(3*n+1))*sin(n) then
m:=m+1:
print([n,m]):
end_if:
end_for:
displays the all known elements of ${\mathcal{X}}$. The output ends with the line $[95151.0,~{}30311]$, which proves Condition (1a) with $n=95151$ and Condition (4a) with ${\rm card}({\mathcal{X}})\geqslant 30311$. ∎
Statement 9.
The set
${\mathcal{X}}=\{k\in\mathbb{N}:{\rm card}([-1,k]\cap{\mathcal{P}}_{n^{2}+1})<10^{10000}\}$
satisfies the conjunction
$\neg(Condition~{}{\tt 1a})\wedge(Condition~{}{\tt 2a})\wedge(Condition~{}{\tt 3a})\wedge(Condition~{}{\tt 4a})\wedge(Condition~{}{\tt 5a})$
Statement 10.
There exists a naturally defined set ${\mathcal{C}}\subseteq\mathbb{N}$ which satisfies the following Conditions (6)-(11). (6) A known and simple algorithm for every $k\in\mathbb{N}$ decides whether or not $k\in{\mathcal{C}}$. (7) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}({\mathcal{C}})=\omega$" is outside ${\mathcal{K}}$. (8) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}(\mathbb{N}\setminus{\mathcal{C}})=\omega$" is outside ${\mathcal{K}}$. (9) It is conjectured, though so far unproven, that ${\mathcal{C}}$ is infinite. (10) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns an integer $n$ satisfying ${\rm card}({\mathcal{C}})<\omega\Rightarrow{\mathcal{C}}\subseteq(-\infty,n]$" is outside ${\mathcal{K}}$. (11) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns an integer $m$ satisfying ${\rm card}(\mathbb{N}\setminus{\mathcal{C}})<\omega\Rightarrow\mathbb{N}\setminus{\mathcal{C}}\subseteq(-\infty,m]$" is outside ${\mathcal{K}}$.
Proof.
Conditions (6)-(11) hold for ${\mathcal{C}}=\{k\in\mathbb{N}:~{}2^{\textstyle 2^{k}}+1~{}is~{}composite\}$. It follows from the following three observations. It is an open problem whether or not there are infinitely many composite numbers of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 159] and [15, p. 74]. It is an open problem whether or not there are infinitely many prime numbers of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 158] and [15, p. 74]. Most mathematicians believe that $12^{\textstyle 2^{k}}+$ 1 is composite for every integer $k\geqslant 5$, see [7, p. 23]. ∎
8 Subsets of $\mathbb{N}$ and their threshold numbers
Definition 4.
We say that an integer $n$ is a threshold number of a set ${\mathcal{X}}\subseteq\mathbb{N}$, if ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\mathcal{X}}\subseteq(-\infty,n]$.
If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is empty or infinite, then any integer $n$ is a threshold number of ${\mathcal{X}}$. If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is non-empty and finite, then the all threshold numbers of ${\mathcal{X}}$ form the set $[\max({\mathcal{X}}),\infty)\cap\mathbb{N}$.
Definition 5.
We say that a non-negative integer $n$ is a weak threshold number of a set ${\mathcal{X}}\subseteq\mathbb{N}$, if ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\rm card}({\mathcal{X}})\leqslant n$.
If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is infinite, then any non-negative integer $n$ is a weak threshold number of ${\mathcal{X}}$. If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is finite, then the all weak threshold numbers of ${\mathcal{X}}$ form the set $[{\rm card}({\mathcal{X}}),\infty)\cap\mathbb{N}$.
Let ${\mathcal{X}}=\{k\in\mathbb{N}:~{}any~{}proof~{}in~{}ZFC~{}of~{}length~{}k~{}or~{}less~{}does~{}not~{}show~{}that~{}0=1\}$.
Lemma 1.
If $n\in\mathbb{N}$ and ${\rm card}({\mathcal{X}})\leqslant n$, then ${\mathcal{X}}\subseteq(-\infty,n-1]$.
Theorem 1.
For every explicitly given $n\in\mathbb{Z}$, if $ZFC$ proves that $n$ is a threshold number of ${\mathcal{X}}$, then $ZFC$ is inconsistent. For every explicitly given $n\in\mathbb{N}$, if $ZFC$ proves that $n$ is a weak threshold number of ${\mathcal{X}}$, then $ZFC$ is inconsistent.
Proof.
If follows from Lemma 1 and the second Gödel incompleteness theorem. ∎
Open Problem 4.
Is there a known (weak) threshold number of ${\mathcal{P}}_{n^{2}+1}$?
Open Problem 5.
Is there a known (weak) threshold number of ${\mathcal{P}}_{n!+1}$?
9 A theorem in the classical computability theory
Let
$E_{n}=\{1=x_{k},~{}x_{i}+x_{j}=x_{k},~{}x_{i}\cdot x_{j}=x_{k}:~{}i,j,k\in\{0,\ldots,n\}\}$
Theorem 2.
([18, p. 118]). There is a limit-computable function $\beta_{1}:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\delta_{1}:\mathbb{N}\to\mathbb{N}$.
We present an alternative proof of Theorem 2. For every $n\in\mathbb{N}$, we define $\beta_{1}(n)$ as the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The function $\beta_{1}:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\delta_{1}:\mathbb{N}\to\mathbb{N}$, see [23]. Flowchart 1 describes a semi-algorithm which computes $\beta_{1}(n)$ in the limit.
Flowchart 1.  A semi-algorithm which computes $\beta_{1}(n)$ in the limit.
Conjecture 2.
([4, pp. 341–342], [12, p. 42], [13]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
10 A statement in computability theory with the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge
${\mathcal{K}}_{f}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known formal theorems. The set ${\mathcal{K}}_{f}$ is time-dependent and publicly available. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}_{f}$. In the statement: It is unknown whether or not there are infinitely many twin primes, the predicate ${\mathcal{K}}_{f}$ appears exclusively outside a formal mathematical statement.
In the statement: It is unknown whether or not there are infinitely many twin primes, the predicate ${\mathcal{K}}_{f}$ appears exclusively outside a formal mathematical statement. Statement 11 is non-trivially true, contains the predicate ${\mathcal{K}}_{f}$, and will be false when someone proves Conjecture 2. Since the function $\beta_{1}$ in Theorem 2 is not computable, Statement 11 does not follow from Theorem 2. Ignoring the epistemic condition in Statement 11, Statement 11 follows from Theorem 2. Statement 11 does not follow from any widely known mathematical theorem. Statement 11 shows that there is a non-trivially true statement in computability (which may be falsified) in which the predicate ${\mathcal{K}}_{f}$ appears exclusively inside the statement.
Statement 11.
There is a limit-computable function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
This is proved in [23]. The term "dominated" in the title of [23] means "eventually dominated". For every $n\in\mathbb{N}$, we define $\beta(n)$ as the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. Flowchart 2 describes a semi-algorithm which computes $\beta(n)$ in the limit.
Flowchart 2.  A semi-algorithm which computes $\beta(n)$ in the limit.
Corollary 3.
The proof of Statement 11 shows that Statement 11 follows from some mathematical theorem and the following conjunction: ((The function $\beta$ is computable) $\not\in{\mathcal{K}}_{f}$) $\wedge$ ((The function $\beta$ is uncomputable) $\not\in{\mathcal{K}}_{f}$).