### v19

# Mathematics with the knowledge predicate ${\mathcal{K}}$ satisfied by every currently known theorem Apoloniusz Tyszka

Abstract
${\mathcal{K}}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known theorems. The set ${\mathcal{K}}$ is time-dependent, publicly available, and contains theorems both from formal and constructive mathematics. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}$. Mathematical statements with known constructive proofs exist in ${\mathcal{K}}$ separately and form the set ${\mathcal{K}}_{\textstyle c}\subseteq{\mathcal{K}}$. We assume that mathematical sets are atemporal entities. They exist formally in $ZFC$ theory although their properties can be time-dependent (when they depend on ${\mathcal{K}}$) or informal. Algorithms always terminate. We explain the distinction between algorithms whose existence is provable in $ZFC$ and constructively defined algorithms which are currently known. By using this distinction, we obtain non-trivially true statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ that belong to constructive and informal mathematics and refer to the current mathematical knowledge on ${\mathcal{X}}$. We present non-trivially true statements in computability with the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge, which do not express the current and formal mathematical knowledge and may be falsified.
Key words and phrases: constructive mathematics, constructively defined algorithms, current mathematical knowledge, informal mathematics, known algorithms, time-dependent notion of truth in mathematics.
2020 Mathematics Subject Classification: 03F65.
1 Introduction
This article in an extended and changed version of [24].
${\mathcal{K}}$ denotes both the knowledge predicate satisfied by every currently known theorem and the finite set of all currently known theorems. The set ${\mathcal{K}}$ is time-dependent, publicly available, and contains theorems both from formal and constructive mathematics. Any theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}$. Mathematical statements with known constructive proofs exist in ${\mathcal{K}}$ separately and form the set ${\mathcal{K}}_{\textstyle c}\subseteq{\mathcal{K}}$. We assume that mathematical sets are atemporal entities. They exist formally in $ZFC$ theory although their properties can be time-dependent (when they depend on ${\mathcal{K}}$) or informal. The true statement "${\mathcal{K}}$ is non-empty" is outside ${\mathcal{K}}$ forever because ${\mathcal{K}}$ is not a formal set.
Paul Cohen proved in $1963$ that the equality $2^{\aleph_{0}}=\aleph_{1}$ is independent of $ZFC$, see [3]. Before $1963$, the statement "There is a constructively defined integer $n\geqslant 1$ such that $2^{\aleph_{0}}=\aleph_{n}$" was outside ${\mathcal{K}}$. Since $1963$, this statement is outside ${\mathcal{K}}$ forever. The true statement "There exists a set ${\mathcal{X}}\subseteq\{1,\ldots,49\}$ such that ${\rm card}({\mathcal{X}})=6$ and ${\mathcal{X}}$ never occurred as the winning six numbers in the Polish Lotto lottery" refers to the current non-mathematical knowledge and is outside ${\mathcal{K}}$ forever.
Algorithms always terminate. Semi-algorithms may not terminate. There is the distinction between existing algorithms (i.e. algorithms whose existence is provable in $ZFC$) and known algorithms (i.e. algorithms whose definition is constructive and currently known), see [2], [17, p. 9]. By using this distinction, we obtain non-trivially true statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ that belong to constructive and informal mathematics and refer to the current mathematical knowledge on ${\mathcal{X}}$. For every such statement $\Phi$, Observations 14 hold.
Observation 1.
The truth of $\Phi$ concerning the current mathematical knowledge is implied by a true statement of the form:
(the conjunction of $i$ conditions of the form $\alpha\in{\mathcal{K}}$) $\wedge$ (the conjunction of $j$ conditions of the form $\beta\not\in{\mathcal{K}}$) $\wedge$
(the conjunction of $k$ conditions of the form $\gamma\in{\mathcal{K}}_{\textstyle c}$) $\wedge$ (the conjunction of $l$ conditions of the form $\delta\not\in{\mathcal{K}}_{\textstyle c}$),
where $i,j,k,l\in\mathbb{N}$ and $\alpha,\beta,\gamma,\delta$ are mathematical statements. In particular, the truth of Statement 6 concerning the current mathematical knowledge guarantees that the set
${\mathcal{X}}=\{n\in\mathbb{N}:~{}the~{}interval~{}[-1,n]~{}contains~{}more~{}than~{}29.5+\frac{\textstyle 11!}{\textstyle 3n+1}\cdot\sin(n)~{}primes~{}of~{}the~{}form~{}k!+1\}$
satisfies the following conjunction:
${\rm(the~{}statement~{}}"the~{}algorithm~{}{\mathcal{A}}_{1}~{}returns~{}the~{}logical~{}value~{}of~{}the~{}statement~{}{\rm card}({\mathcal{X}})=\omega"{\rm~{}is~{}outside~{}{\mathcal{K}})\wedge\ldots\wedge}$
${\rm(the~{}statement~{}}"the~{}algorithm~{}{\mathcal{A}}_{p}~{}returns~{}the~{}logical~{}value~{}of~{}the~{}statement~{}{card}({\mathcal{X}})=\omega"{\rm~{}is~{}outside~{}{\mathcal{K}}),}$
where ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{p}$ are the all known algorithms with no input. Statement 6 will be false when someone proves that there is a constructively defined algorithm ${\mathcal{A}}$ with no input that returns the logical value of the statement ${\rm card}({\mathcal{X}})=\omega$.
Observation 2.
The proof of $\Phi$ uses mathematical theorems. For example, the proof of Statement 6 uses the fact that the set ${\mathcal{X}}$ from Observation 1 satisfies ${\rm card}({\mathcal{X}})<\omega\Rightarrow{\mathcal{X}}\subseteq(-\infty,501893]$.
Observation 3.
The proof of $\Phi$ does not use the assumption that we can list the all elements of ${\mathcal{K}}$.
Observation 4.
The proof of $\Phi$ does not use the assumption that we can list the all elements of ${\mathcal{K}}_{\textstyle c}$.
Observation 5 is known from the beginning of computability theory and shows that the predicate ${\mathcal{K}}$ increases intuitive mathematics.
Observation 5.
Church’s thesis is based on the fact that the all currently known intuitively computable functions are computable.
In Observation 6, the predicate ${\mathcal{K}}$ trivially increases constructive mathematics.
Observation 6.
The largest known prime number has the form $12^{\textstyle n}-$ 1.
2 Time-dependent notion of truth in constructive mathematics
Below is the English summary of [22] available at the internet address of [22].
The basic philosophical idea of intuitionism is that mathematical entities exist only as mental constructions and that the notion of truth of a proposition should be equated with its verification or the existence of proof. However different intuitionists explained the existence of a proof in fundamentally different ways. There seem to be two main alternatives: the actual and potential existence of a proof. The second proposal is also understood in two alternative ways: as knowledge of a method of construction of a proof or as knowledge-independent and tenseless existence of a proof. This paper is a presentation and analysis of these alternatives.
In constructive mathematics ([16]) and the traditional Brouwerian intuitionism ([11, p. 135]), the truth of a mathematical statement means that we know a constructive proof. Therefore, the truth of a mathematical statement depends on time, where the statement is formally stated in the classical mathematics without the predicate ${\mathcal{K}}$.
In this article, mathematical statements on decidable sets ${\mathcal{X}}\subseteq\mathbb{N}$ refer to time because they refer to the current mathematical knowledge on ${\mathcal{X}}$. They cannot be formally stated in the classical mathematics without the predicate ${\mathcal{K}}$ and their logical values may change in time.
In Martin-Löf’s terminology ([10, p. 142]), every currently known theorem is actually true whereas every theorem (known or unknown) is potentially true. Actual truth is knowledge dependent and tensed. Potential truth is knowledge independent and tenseless.
3 Basic definitions and examples
Definition 1 applies to sets ${\mathcal{X}}\subseteq\mathbb{N}$ whose infiniteness is false or unproven.
Definition 1.
[…]
The set of known elements of ${\mathcal{X}}$ is finite and time-dependent, so cannot be defined in the formal language of classical mathematics. Let $t$ denote the largest twin prime that is smaller than ((((((((9!)!)!)!)!)!)!)!)!. The number $t$ is an unknown element of the set of twin primes.
Definition 2.
Conditions (1)-(5) concern sets ${\mathcal{X}}\subseteq\mathbb{N}$. (1) A known algorithm with no input returns an integer $n$ satisfying ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\mathcal{X}}\subseteq(-\infty,n]$. (2) A known algorithm for every $k\in\mathbb{N}$ decides whether or not $k\in{\mathcal{X}}$. (3) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}({\mathcal{X}})=\omega$" is outside ${\mathcal{K}}$. (4) There are many elements of ${\mathcal{X}}$ and it is conjectured, though so far unproven, that ${\mathcal{X}}$ is infinite. (5) ${\mathcal{X}}$ is naturally defined. The infiniteness of ${\mathcal{X}}$ is false or unproven. ${\mathcal{X}}$ has the simplest definition among known sets ${\mathcal{Y}}\subseteq\mathbb{N}$ with the same set of known elements.
Condition (3) implies that no known proof shows the finiteness/infiniteness of ${\mathcal{X}}$. No known set ${\mathcal{X}}\subseteq\mathbb{N}$ satisfies Conditions (1)-(4) and is widely known in number theory or naturally defined, where this term has only informal meaning.
Let []delimited-[][$\cdot$][ ⋅ ] denote the integer part function.
Example 1.
The set
${\mathcal{X}}=\left\{\begin{array}[]{cl}\mathbb{N},&{\rm if~{}}[\frac{((((((((9!)!)!)!)!)!)!)!)!}{\pi}]~{}is~{}odd\\ \emptyset,&otherwise\end{array}\right.$
does not satisfy Condition (3) because we know an algorithm with no input that computes $[\frac{((((((((9!)!)!)!)!)!)!)!)!}{\pi}]$. The set of known elements of ${\mathcal{X}}$ is empty. Hence, Condition (5) fails for ${\mathcal{X}}$.
Example 2.
([2][17, p. 9]). The function
$\mathbb{N}\ni n\stackrel{{\scriptstyle\textstyle h}}{{\longrightarrow}}\left\{\begin{array}[]{rl}1,&if~{}the~{}decimal~{}expansion~{}of~{}\pi~{}contains~{}n~{}consecutive~{}zeros\\ 0,&otherwise\end{array}\right.$
is computable because $h=\mathbb{N}\times\{1\}$ or there exists $k\in\mathbb{N}$ such that
$h=(\{0,\ldots,k\}\times\{1\})\cup(\{k+1,~{}k+2,~{}k+3,~{}\ldots\}\times\{0\})$
No known algorithm computes the function $h$.
Example 3.
The set
${\mathcal{X}}=\left\{\begin{array}[]{rl}\mathbb{N},&if~{}the~{}continuum~{}hypothesis~{}holds\\ \emptyset,&otherwise\end{array}\right.$
is decidable. This ${\mathcal{X}}$ satisfies Conditions (1) and (3) and does not satisfy Conditions (2), (4), and (5). These facts will hold forever.
4 A consequence of the physical limits of computation
Statement 1.
No set ${\mathcal{X}}\subseteq\mathbb{N}$ will satisfy Conditions (1)-(4) forever, if for every algorithm with no input, at some future day, a computer will be able to execute this algorithm in $1$ second or less.
Proof.
The proof goes by contradiction. We fix an integer $n$ that satisfies Condition (1). Since Conditions (1)-(3) will hold forever, the semi-algorithm in Figure 1 never terminates and sequentially prints the following sentences:
$n+1\not\in{\mathcal{X}},~{}n+2\not\in{\mathcal{X}},~{}n+3\not\in{\mathcal{X}},~{}\ldots$ (T)
Figure 1.   Semi-algorithm that terminates if and only if ${\mathcal{X}}$ is infinite.
The sentences from the sequence (T) and our assumption imply that for every integer $m>n$ computed by a known algorithm, at some future day, a computer will be able to confirm in $1$ second or less that $(n,m]\cap{\mathcal{X}}=\emptyset$. Thus, at some future day, numerical evidence will support the conjecture that the set ${\mathcal{X}}$ is finite, contrary to the conjecture in Condition (4). ∎
The physical limits of computation ([9]) disprove the assumption of Statement 1.
5 Statements which refer to Conditions (1)-(5)
Edmund Landau’s conjecture states that the set ${\mathcal{P}}_{n^{2}+1}$ of primes of the form $n^{2}+1$ is infinite, see [20], [21], [27].
Statement 2.
Condition (1) remains unproven for ${\mathcal{X}}={\mathcal{P}}_{n^{2}+1}$.
Proof.
For every set ${\mathcal{X}}\subseteq\mathbb{N}$, there exists an algorithm ${\tt Alg}({\mathcal{X}})$ with no input that returns
$n=\left\{\begin{array}[]{rl}0,&{\rm if~{}}{\rm card}({\mathcal{X}})\in\{0,\omega\}\\ \max({\mathcal{X}}),&{\rm otherwise}\end{array}\right.$
This $n$ satisfies the implication in Condition (1), but the algorithm ${\tt Alg}({\mathcal{P}}_{n^{2}+1})$ is unknown because its definition is ineffective. ∎
Statement 3.
The statement
$\exists n\in\mathbb{N}~{}({\rm card}({\mathcal{P}}_{n^{2}+1})<\omega\Rightarrow{\mathcal{P}}_{n^{2}+1}\subseteq[2,n+3])$
remains unproven in $ZFC$ and classical logic without the law of excluded middle.
Statement 4.
The set
${\mathcal{X}}=\{k\in\mathbb{N}:~{}{\rm card}({\mathcal{P}}_{n^{2}+1}\cap[-1,k])>10^{\textstyle 10^{\textstyle 10}}\}\cup\{k\in{\mathcal{P}}_{n^{2}+1}:~{}{\rm card}({\mathcal{P}}_{n^{2}+1}\cap[-1,k])\leqslant 10^{\textstyle 10^{\textstyle 10}}\}$
satisfies Conditions (2)-(4). Condition (1) fails for ${\mathcal{X}}$, ${\rm card}({\mathcal{X}})<\omega\Rightarrow{\rm card}({\mathcal{X}})\leqslant 10^{\textstyle 10^{\textstyle 10}}$.
Proof.
Since ${\rm card}({\mathcal{P}}_{n^{2}+1}\cap[2,10^{28}))=2199894223892$ ([21]) and the inequality ${\rm card}(P_{n^{2}+1})\geqslant$ $10^{\textstyle 10^{\textstyle 10}}$ remains unproven, Conditions (3) and (4) hold. ∎
For a non-negative integer $n$, let $\theta(n)$ denote the largest integer divisor of $10^{\textstyle 10^{\textstyle 10}}$ smaller than $n$. Let $\kappa:\mathbb{N}\to\mathbb{N}$ be defined by setting $\kappa(n)$ to be the exponent of $2$ in the prime factorization of $n+1$.
Statement 5.
([25, p. 250]). The set
${\mathcal{X}}=\{n\in\mathbb{N}:(\theta(n)+\kappa(n))^{2}+1~{}is~{}prime\}$
satisfies Conditions (1)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined. Condition (1) holds with $n=10^{\textstyle 10^{\textstyle 10}}$.
Let ${\mathcal{P}}_{n!+1}$ denote the set of primes of the form $n!+1$.
Conjecture 1.
([1, p. 443], [6]). The set ${\mathcal{P}}_{n!+1}$ is infinite.
For a non-negative integer $n$, let $\rho(n)$ denote $29.5+\frac{\textstyle 11!}{\textstyle 3n+1}\cdot\sin(n)$.
Statement 6.
The set
${\mathcal{X}}=\{n\in\mathbb{N}:the~{}interval~{}[-1,n]~{}contains~{}more~{}than~{}\rho(n)~{}elements~{}of~{}{\mathcal{P}}_{n!+1}\}$
satisfies Conditions (1)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined. $501893\in{\mathcal{X}}$. Condition (1) holds with $n=501893$. ${\rm card}({\mathcal{X}}\cap[0,501893])=159827$. ${\mathcal{X}}\cap[501894,\infty)=$$\{n\in\mathbb{N}:$ $the$ $interval$ $[-1,n]$ $contains$ $at$ $least$ $30$ $elements$ $of$ ${\mathcal{P}}_{n!+1}\}$.
Proof.
For every integer $n\geqslant 11!$, $30$ is the smallest integer greater than $\rho(n)$. By this, if $n\in{\mathcal{X}}\cap[11!,\infty)$, then $n+1,~{}n+2,~{}n+3,~{}\ldots~{}\in{\mathcal{X}}$. Hence, Condition (1) holds with $n=11!-1$. Since the inequality ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 30$ remains unproven, Condition (3) holds. The interval $[-1,11!-1]$ contains exactly three primes of the form $k!+1$: $1!+1$, $2!+1$, $3!+1$. For every integer $n>503000$, the inequality $\rho(n)>3$ holds. Therefore, the execution of the following MuPAD code
m:=0:
for n from 0.0 to 503000.0 do
if n<1!+1 then r:=0 end_if:
if n>=1!+1 and n<2!+1 then r:=1 end_if:
if n>=2!+1 and n<3!+1 then r:=2 end_if:
if n>=3!+1 then r:=3 end_if:
if r>29.5+(11!/(3*n+1))*sin(n) then
m:=m+1:
print([n,m]):
end_if:
end_for:
displays the all known elements of ${\mathcal{X}}$. The output ends with the line $[501893.0,~{}159827]$, which proves Condition (1) with $n=501893$ and Condition (4) with ${\rm card}({\mathcal{X}})\geqslant 159827$. ∎
T. Nagell proved in [14] (cf. [19, p. 104]) that the equation $x^{2}-17=y^{3}$ has exactly $16$ integer solutions, namely (±$(\pm 3,-2)$, (±$(\pm 4,-1)$, (±$(\pm 5,2)$, (±$(\pm 9,4)$, (±$(\pm 23,8)$, (±$(\pm 282,43)$, (±$(\pm 375,52)$, (±$(\pm 378661,5234)$. The set
$\bigcup_{\stackrel{{\scriptstyle\textstyle(x,y)\in\mathbb{Z}\times\mathbb{Z}}}{{\textstyle(x^{2}-y^{3}-17)\cdot(y^{2}-x^{3}-17)=0}}}\{(x+8)^{8}\}$
has exactly $23$ elements. Among them, there are $14$ integers from the interval $[1,2199894223892]$. Let ${\mathcal{W}}$ denote the set
$\bigcup_{\stackrel{{\scriptstyle\textstyle(x,y)\in\mathbb{Z}\times\mathbb{Z}}}{{\textstyle(x^{2}-y^{3}-17)\cdot(y^{2}-x^{3}-17)=0}}}\{k\in\mathbb{N}:~{}k~{}is~{}the~{}(x+8)^{8}-th~{}element~{}of~{}{\mathcal{P}}_{n^{2}+1}\}$
From [21], it is known that ${\rm card}({\mathcal{P}}_{n^{2}+1}\cap[2,10^{28}))=2199894223892$. Hence, ${\rm card}({\mathcal{W}}\cap[2,10^{28}))=14$ and $14$ elements of ${\mathcal{W}}$ can be practically computed. The inequality ${\rm card}({\mathcal{P}}_{n^{2}+1})\geqslant(378661+8)^{8}$ remains unproven. The last two sentences and Statement 6 imply the following corollary.
Corollary 1.
If we add ${\mathcal{W}}$ to ${\mathcal{X}}$, then the following statements hold: Condition (1) fails for ${\mathcal{X}}$, $159827+14\leqslant{\rm card}({\mathcal{X}})$, the above lower bound is currently the best known, ${\rm card}({\mathcal{X}})<\omega\Rightarrow{\rm card}({\mathcal{X}})\leqslant 159827+23$, the above upper bound is currently the best known, ${\mathcal{X}}$ satisfies Conditions (2)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined.
Corollary 2.
Since the inequality ${\rm card}({\mathcal{P}}_{n^{2}+1})>9^{\textstyle 999}$ remains unproven and $10^{\textstyle 953}<9^{\textstyle 999}<10^{\textstyle 954}$, analogical statements hold when we add to ${\mathcal{X}}$ the set
$\bigcup_{\textstyle i\in\mathbb{N}}\left\{k\in\mathbb{N}:~{}k-501894~{}is~{}the~{}\left(\left[\frac{9^{\textstyle 999}}{10^{\textstyle i}}\right]+1\right)-th~{}element~{}of~{}{\mathcal{P}}_{n^{2}+1}\right\}$
which has at most $955$ elements.
For a non-negative integer $i$, let $d(i)$ denote the smallest prime divisor of $\left[31+\frac{\textstyle 10^{\textstyle 6}}{\textstyle i+1}\right]$.
Statement 7.
The set
${\mathcal{X}}=\bigcup_{\textstyle i\in\mathbb{N}}\left\{k^{\textstyle i}:~{}k~{}is~{}the~{}d(i)-th~{}element~{}of~{}{\mathcal{P}}_{n!+1}\right\}$
satisfies Conditions (2)-(5) except the requirement that ${\mathcal{X}}$ is naturally defined. Condition (1) fails for ${\mathcal{X}}$. ${\rm card}({\mathcal{X}})\geqslant 946732$. If ${\rm card}({\mathcal{P}}_{n!+1})\leqslant 28$, then ${\rm card}({\mathcal{X}})=946732$. If $29\leqslant{\rm card}({\mathcal{P}}_{n!+1})\leqslant 30$, then ${\rm card}({\mathcal{X}})=946745$. If ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 31$, then ${\rm card}({\mathcal{X}})=\omega$.
Proof.
The inequality ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 23$ holds, see [5]. The inequality ${\rm card}({\mathcal{P}}_{n!+1})\geqslant 29$ remains unproven. The execution of the following MuPAD code
[m,n]:=[0,0]:
for i from 0 to 10^6-1 do
A:=numlib::primedivisors(floor(31+(10^6/(i+1)))):
if A[1]<=23 then m:=m+1 end_if:
if A[1]<=29 then n:=n+1 end_if:
end_for:
print([m,n]):
displays $[946732,946745]$. The last claim of Statement 7 holds because $d\left(\left[31+\frac{\textstyle 10^{\textstyle 6}}{\textstyle i+1}\right]\right)=31$ for every integer $i\geqslant 10^{\textstyle 6}$. Condition (1) fails for ${\mathcal{X}}$ because we cannot rule out the possibility that $29\leqslant{\rm card}({\mathcal{P}}_{n!+1})\leqslant 30$. ∎
6 Satisfiable conjunctions which consist of Conditions (1)-(5) and their negations
Open Problem 1.
Is there a set ${\mathcal{X}}\subseteq\mathbb{N}$ which satisfies Conditions (1)-(5)?
Open Problem 1 asks about the existence of a year $t\geqslant 2024$ in which the conjunction
$(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})$
will hold for some ${\mathcal{X}}\subseteq\mathbb{N}$. For every year $t\geqslant 2024$ and for every $i\in\{1,2,3\}$, a positive solution to Open Problem $i$ in the year $t$ may change in the future. Currently, the answers to Open Problems 15 are negative.
The set ${\mathcal{X}}={\mathcal{P}}_{n^{2}+1}$ satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})$
The set ${\mathcal{X}}=\{0,\ldots,10^{6}\}\cup{\mathcal{P}}_{n^{2}+1}$ satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge\neg(Condition~{}{\tt 5})$
Let $f(1)=10^{6}$, and let $f(n+1)=f(n)^{f(n)}$ for every positive integer $n$. The set
${\mathcal{X}}=\left\{\begin{array}[]{l}\mathbb{N},~{}if~{}2^{\textstyle 2^{\textstyle f(9^{9})}}+1~{}is~{}composite\\ \{0,\ldots,10^{6}\},~{}otherwise\end{array}\right.$
satisfies the conjunction
$(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge\neg(Condition~{}{\tt 5})$
Open Problem 2.
Is there a set ${\mathcal{X}}\subseteq\mathbb{N}$ that satisfies the conjunction
$(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})?$
The numbers $12^{\textstyle 2^{k}}+$ 1 are prime for $k\in\{0,1,2,3,4\}$. It is open whether or not there are infinitely many primes of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 158] and [15, p. 74]. It is open whether or not there are infinitely many composite numbers of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 159] and [15, p. 74]. Most mathematicians believe that $12^{\textstyle 2^{k}}+$ 1 is composite for every integer $k\geqslant 5$, see [7, p. 23]. The set
${\mathcal{X}}=\left\{\begin{array}[]{l}\mathbb{N},~{}if~{}2^{\textstyle 2^{\textstyle f(9^{9})}}+1~{}is~{}composite\\ \{0,\ldots,10^{6}\}\cup\\ \{n\in\mathbb{N}:n~{}is~{}the~{}sixth~{}prime~{}number~{}of~{}the~{}form~{}2^{\textstyle 2^{\textstyle k}}+1\},~{}otherwise\end{array}\right.$
satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge\neg(Condition~{}{\tt 5})$
Open Problem 3.
Is there a set ${\mathcal{X}}\subseteq\mathbb{N}$ that satisfies the conjunction
$\neg(Condition~{}{\tt 1})\wedge(Condition~{}{\tt 2})\wedge\neg(Condition~{}{\tt 3})\wedge(Condition~{}{\tt 4})\wedge(Condition~{}{\tt 5})?$
It is possible, although very doubtful, that at some future day, the set ${\mathcal{X}}={\mathcal{P}}_{n^{2}+1}$ will solve Open Problem 2. The same is true for Open Problem 3. It is possible, although very doubtful, that at some future day, the set ${\mathcal{X}}=\{k\in\mathbb{N}:2^{\textstyle 2^{k}}+1~{}is~{}composite\}$ will solve Open Problem 1. The same is true for Open Problems 2 and 3.
Table 1 shows satisfiable conjunctions of the form
#$\#{\rm(Condition}~{}{\tt 1}{\rm)}\wedge{\rm(Condition}~{}{\tt 2}{\rm)}\wedge\#{\rm(Condition}~{}{\tt 3}{\rm)}\wedge{\rm(Condition}~{}{\tt 4}{\rm)}\wedge\#{\rm(Condition}~{}{\tt 5}{\rm)}$
where ##\## denotes the negation $\neg$ or the absence of any symbol.
Table 1.   Five satisfiable conjunctions.
7 Statements which refer to Conditions (1a)-(5a) and (6)-(11)
Definition 3.
Conditions (1a)-(5a) concern sets ${\mathcal{X}}\subseteq\mathbb{N}$. (1a) A known algorithm with no input returns a positive integer $n$ satisfying ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\mathcal{X}}\subseteq(-\infty,n]$. (2a) A known algorithm for every $k\in\mathbb{N}$ decides whether or not $k\in{\mathcal{X}}$. (3a) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}({\mathcal{X}})<\omega$" is outside ${\mathcal{K}}$. (4a) There are many elements of ${\mathcal{X}}$ and it is conjectured, though so far unproven, that ${\mathcal{X}}$ is finite. (5a) ${\mathcal{X}}$ is naturally defined. The finiteness of ${\mathcal{X}}$ is false or unproven. ${\mathcal{X}}$ has the simplest definition among known sets ${\mathcal{Y}}\subseteq\mathbb{N}$ with the same set of known elements.
Statement 8.
The set
${\mathcal{X}}=\Bigl{\{}n\in\mathbb{N}:~{}the~{}interval~{}[-1,n]~{}contains~{}more~{}than~{}6.5+\frac{10^{6}}{3n+1}\cdot\sin(n)~{}squares~{}of~{}the~{}form~{}k!+1\Bigr{\}}$
satisfies Conditions (1a)-(5a) except the requirement that ${\mathcal{X}}$ is naturally defined. $95151\in{\mathcal{X}}$. Condition (1a) holds with $n=95151$. ${\rm card}({\mathcal{X}}\cap[0,95151])=30311$. ${\mathcal{X}}\cap[95152,\infty)=$$\{n\in\mathbb{N}:$ $the$ $interval$ $[-1,n]$ $contains$ $at$ $least$ $7$ $squares$ $of$ $the$ $form$ $k!+1\}$.
Proof.
For every integer $n>10^{6}$, $7$ is the smallest integer greater than $6.5+\frac{10^{6}}{3n+1}\cdot\sin(n)$. By this, if $n\in{\mathcal{X}}\cap(10^{6},\infty)$, then $n+1,~{}n+2,~{}n+3,~{}\ldots~{}\in{\mathcal{X}}$. Hence, Condition (1a) holds with $n=10^{6}$. It is conjectured that $k!+1$ is a square only for $k\in\{4,5,7\}$, see [26, p. 297]. Hence, the inequality ${\rm card}(\{k\in\mathbb{N}\setminus\{0\}:k!+1~{}is~{}a~{}square\})>3$ remains unproven. Since $3<7$, Condition (3a) holds. The interval $[-1,10^{6}]$ contains exactly three squares of the form $k!+1$: $4!+1$, $5!+1$, $7!+1$. Therefore, the execution of the following MuPAD code
m:=0:
for n from 0.0 to 1000000.0 do
if n<25 then r:=0 end_if:
if n>=25 and n<121 then r:=1 end_if:
if n>=121 and n<5041 then r:=2 end_if:
if n>=5041 then r:=3 end_if:
if r>6.5+(1000000/(3*n+1))*sin(n) then
m:=m+1:
print([n,m]):
end_if:
end_for:
displays the all known elements of ${\mathcal{X}}$. The output ends with the line $[95151.0,~{}30311]$, which proves Condition (1a) with $n=95151$ and Condition (4a) with ${\rm card}({\mathcal{X}})\geqslant 30311$. ∎
Statement 9.
The set
${\mathcal{X}}=\{k\in\mathbb{N}:{\rm card}([-1,k]\cap{\mathcal{P}}_{n^{2}+1})<10^{10000}\}$
satisfies the conjunction
$\neg(Condition~{}{\tt 1a})\wedge(Condition~{}{\tt 2a})\wedge(Condition~{}{\tt 3a})\wedge(Condition~{}{\tt 4a})\wedge(Condition~{}{\tt 5a})$
Statement 10.
There exists a naturally defined set ${\mathcal{C}}\subseteq\mathbb{N}$ which satisfies the following Conditions (6)-(11). (6) A known and simple algorithm for every $k\in\mathbb{N}$ decides whether or not $k\in{\mathcal{C}}$. (7) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}({\mathcal{C}})=\omega$" is outside ${\mathcal{K}}$. (8) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns the logical value of the statement ${\rm card}(\mathbb{N}\setminus{\mathcal{C}})=\omega$" is outside ${\mathcal{K}}$. (9) It is conjectured, though so far unproven, that ${\mathcal{C}}$ is infinite. (10) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns an integer $n$ satisfying ${\rm card}({\mathcal{C}})<\omega\Rightarrow{\mathcal{C}}\subseteq(-\infty,n]$" is outside ${\mathcal{K}}$. (11) For every known algorithm ${\mathcal{A}}$ with no input, the statement "${\mathcal{A}}$ returns an integer $m$ satisfying ${\rm card}(\mathbb{N}\setminus{\mathcal{C}})<\omega\Rightarrow\mathbb{N}\setminus{\mathcal{C}}\subseteq(-\infty,m]$" is outside ${\mathcal{K}}$.
Proof.
Conditions (6)-(11) hold for ${\mathcal{C}}=\{k\in\mathbb{N}:~{}2^{\textstyle 2^{k}}+1~{}is~{}composite\}$. It follows from the following three observations. It is an open problem whether or not there are infinitely many composite numbers of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 159] and [15, p. 74]. It is an open problem whether or not there are infinitely many prime numbers of the form $12^{\textstyle 2^{k}}+$ 1, see [8, p. 158] and [15, p. 74]. Most mathematicians believe that $12^{\textstyle 2^{k}}+$ 1 is composite for every integer $k\geqslant 5$, see [7, p. 23]. ∎
8 Subsets of $\mathbb{N}$ and their threshold numbers
Definition 4.
We say that an integer $n$ is a threshold number of a set ${\mathcal{X}}\subseteq\mathbb{N}$, if ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\mathcal{X}}\subseteq(-\infty,n]$.
If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is empty or infinite, then any integer $n$ is a threshold number of ${\mathcal{X}}$. If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is non-empty and finite, then the all threshold numbers of ${\mathcal{X}}$ form the set $[\max({\mathcal{X}}),\infty)\cap\mathbb{N}$.
Definition 5.
We say that a non-negative integer $n$ is a weak threshold number of a set ${\mathcal{X}}\subseteq\mathbb{N}$, if ${\rm card}({\mathcal{X}})<\omega\Rightarrow$ ${\rm card}({\mathcal{X}})\leqslant n$.
If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is infinite, then any non-negative integer $n$ is a weak threshold number of ${\mathcal{X}}$. If a set ${\mathcal{X}}\subseteq\mathbb{N}$ is finite, then the all weak threshold numbers of ${\mathcal{X}}$ form the set $[{\rm card}({\mathcal{X}}),\infty)\cap\mathbb{N}$.
Let ${\mathcal{X}}=\{k\in\mathbb{N}:~{}any~{}proof~{}in~{}ZFC~{}of~{}length~{}k~{}or~{}less~{}does~{}not~{}show~{}that~{}0=1\}$.
Lemma 1.
If $n\in\mathbb{N}$ and ${\rm card}({\mathcal{X}})\leqslant n$, then ${\mathcal{X}}\subseteq(-\infty,n-1]$.
Theorem 1.
For every explicitly given $n\in\mathbb{Z}$, if $ZFC$ proves that $n$ is a threshold number of ${\mathcal{X}}$, then $ZFC$ is inconsistent. For every explicitly given $n\in\mathbb{N}$, if $ZFC$ proves that $n$ is a weak threshold number of ${\mathcal{X}}$, then $ZFC$ is inconsistent.
Proof.
If follows from Lemma 1 and the second Gödel incompleteness theorem. ∎
Open Problem 4.
Is there a known (weak) threshold number of ${\mathcal{P}}_{n^{2}+1}$?
Open Problem 5.
Is there a known (weak) threshold number of ${\mathcal{P}}_{n!+1}$?
9 Classical computability theory
Let
$E_{n}=\{1=x_{k},~{}x_{i}+x_{j}=x_{k},~{}x_{i}\cdot x_{j}=x_{k}:~{}i,j,k\in\{0,\ldots,n\}\}$
Theorem 2.
([18, p. 118]). There exists a limit-computable function $\beta_{1}:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\delta_{1}:\mathbb{N}\to\mathbb{N}$.
We present an alternative proof of Theorem 2. For every $n\in\mathbb{N}$, we define $\beta_{1}(n)$ as the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The function $\beta_{1}:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\delta_{1}:\mathbb{N}\to\mathbb{N}$, see [23]. Flowchart 1 describes a semi-algorithm which computes $\beta_{1}(n)$ in the limit.
Flowchart 1
A semi-algorithm which computes $\beta_{1}(n)$ in the limit
Conjecture 2.
([4, pp. 341–342], [12, p. 42], [13]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
For every $n\in\mathbb{N}$, we define $\beta(n)$ as the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
Theorem 3.
The function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
This is proved in [23]. The term "dominated" in the title of [23] means "eventually dominated". Flowchart 2 describes a semi-algorithm which computes $\beta(n)$ in the limit.
Flowchart 2
A semi-algorithm which computes $\beta(n)$ in the limit
10 Non-trivially true statements in computability with the predicate ${\mathcal{K}}_{f}$ of the current and formal mathematical knowledge, which do not express the current and formal mathematical knowledge and may be falsified
${\mathcal{K}}_{f}$ denotes both the knowledge predicate satisfied by every currently known formal theorem and the finite set of all currently known formal theorems. The set ${\mathcal{K}}_{f}$ is time-dependent and publicly available. Any formal theorem of any mathematician from past or present forever belongs to ${\mathcal{K}}_{f}$. The statement $\neg{\mathcal{K}}_{f}(P\neq NP)\wedge\neg{\mathcal{K}}_{f}(P=NP)$ is true and does not justify the title of the section.
Statement 11.
There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
It follows from Theorem 3 by taking $f=\beta$ and the following conjunction:
$(({\rm The~{}function~{}}\beta{\rm~{}is~{}computable)~{}}\not\in{~{}\mathcal{K}}_{f})~{}\wedge~{}(({\rm The~{}function~{}}\beta{\rm~{}is~{}uncomputable)~{}}\not\in~{}{\mathcal{K}}_{f})$
Statement 12.
For every computable function $g:\mathbb{N}\to\mathbb{N}$, there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ of unknown computability such that $f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation and $f(n)>g(n)$ for every $n\in\mathbb{N}$.
Proof.
It follows from Theorem 3 by taking $f=g+1+\beta$ and the following conjunction:
$(({\rm The~{}function~{}}\beta{\rm~{}is~{}computable)~{}}\not\in{~{}\mathcal{K}}_{f})~{}\wedge~{}(({\rm The~{}function~{}}\beta{\rm~{}is~{}uncomputable)~{}}\not\in~{}{\mathcal{K}}_{f})$
Both Statement 11 and Statement 12 justify the title of the section. Statements 11 and 12 are non-trivially true, contain the predicate ${\mathcal{K}}_{f}$, and will be false when someone proves Conjecture 2. Since the function $\beta_{1}$ in Theorem 2 is not computable, Statements 11 and 12 do not follow from Theorem 2. Ignoring the epistemic condition in Statements 11 and 12, they follow from Theorem 2 by taking $f=\beta_{1}$ or $f=g+1+\beta_{1}$.